OpenStudy (anonymous):

simplfiy square root of 98a^17

OpenStudy (anonymous):

i know it breaks down to 49 times 2a^16 times a but then i get lost

OpenStudy (anonymous):

\[7\sqrt{2}\times \sqrt{a ^{17}}\]

myininaya (myininaya):

17/2=8.5 so a^(17/2)=a^8*a^(1/2)

myininaya (myininaya):

so your final answer is 7a^8*sqrt(2a)

OpenStudy (anonymous):

you can simplify the \[\sqrt{a ^{17}}\] to \[a^{8}\sqrt{a}\]

myininaya (myininaya):

very good!

OpenStudy (anonymous):

do you happen to know what the cubed square root of -64 is